The table below shows the annual incomes of the only residents of two countries A and B. Draw each countrys Lorenz curve and determine which country has a more unequal income distribution. 8. Person Incomes in Country A (S)20 25 1000 30 50 40 50 2000 60 100
Answer #1

To construct Lorenze curve from a group of 5 people, order income of people from lowest to highest (as in the below tables):

Country A:

Population % Cumulative % of Population Income A Cumulative Income A Cumulative % Income A
0.00% 0% 0 0 0 0.00%
(1/5)*100 = 20% 20% 20 20 (20/1125)*100 = 1.78%
(1/5)*100 = 20% 40% 25 45 (45/1125)*100 = 4.00%
(1/5)*100 = 20% 60% 30 75 (75/1125)*100 = 6.67%
(1/5)*100 = 20% 80% 50 125 (125/1125)*100 = 11.11%
(1/5)*100 = 20% 100% 1000 1125 (1125/1125)*100 = 100.00%

Country B:

Population % Cumulative % of Population Income B Cumulative Income B Cumulative % Income B
0.00% 0% 0 0 0 0.00%
(1/5)*100 = 20% 20% 40 40 (40/2250)*100 = 1.78%
(1/5)*100 = 20% 40% 50 90 (90/2250)*100 = 4.00%
(1/5)*100 = 20% 80% 60 150 (150/2250)*100 = 6.67%
(1/5)*100 = 20% 100% 100 250 (250/2250)*100 = 11.11%
(1/5)*100 = 20% 60% 2000 2250 (2250/2250)*100 = 100.00%

As can be seen in the graph, Lorenz curves for both countries (red and green) coincide which means that both countries have same level of unequal income distribution.120.00% 100.00% 80.00% Cumulative % Income A Cumulative % Income B Line of Perfect Equality 60.00% 40.00% 20.00% 0.00% + 0.00